At one point during the synthesis of nirosyl bromide, the following concentration were measured [NO] = 0.050 M. [Br2] = 0.100 M, and [NOBr] = 0.010 M. Calculate for the value of Q during this point of the synthesis.
1
Expert's answer
2020-11-02T08:58:46-0500

2NO + Br2 → 2NOBr

The reaction quotient (Q) gives the same ratio the equilibrium expression gives, but for a system that is not at equilibrium.

"Q = \frac{[NOBr]^2}{[NO]^2 \times [Br_2]}"

"Q = \frac{(0.01)^2}{(0.05)^2 \times (0.1)} = 0.4"

Answer: 0.4